OpenStudy (anonymous):

A very small object with mass 6.90 × 10^−9 kg and positive charge 8.30 × 10^−9 C is projected directly toward a very large insulating sheet of positive charge that has uniform surface charge density 5.90 x 10^-8 C/m^2. The object is initially 0.530 m from the sheet. What initial speed must the object have in order for its closest distance of approach to the sheet to be 0.170 m ?

OpenStudy (anonymous):

help guys?

OpenStudy (anonymous):

I'm new to Gauss' law, so this will probably be incorrect. 

OpenStudy (anonymous):

\[\phi =\frac{\sigma A}{\epsilon_0}=\int\limits_{}^{}E(x)dA\]

OpenStudy (anonymous):

E is not a function of x, so This simplifies to:\[\frac{\sigma A}{\epsilon_0}=E \int\limits_{}^{} dA=EA\]

OpenStudy (anonymous):

Sorry, that should have been a half in front of all the fluxes (multiply all epsilon 0s by 2), as only half of the flux goes through the top.

OpenStudy (anonymous):

\[E=\frac{\sigma}{2 \epsilon_0}\]

OpenStudy (anonymous):

\[F=Eq\]

OpenStudy (anonymous):

Okay, let's assume that's true, and the rest is easy.

OpenStudy (anonymous):

\[F=\frac{\sigma q}{2\epsilon_0}=ma\]

OpenStudy (anonymous):

\[6.9*10^{-9}a=\frac{(8.3*10^{-9})(5.9*10^{-8})}{2(8.85*10^{-12})}\]

OpenStudy (anonymous):

Then I use this to solve for a?

OpenStudy (anonymous):

\[a \approx 4*10^3 \pm0.006\]

OpenStudy (anonymous):

Do you understand what I've done so far?

OpenStudy (anonymous):

Anyway, so we have\[a=-4*10^3\]\[s(0)=0.53\] At \[s=0.17\] We want\[v=0\]

OpenStudy (anonymous):

I've got my co-ordinate system muddled, make that a positive acceleration.

OpenStudy (anonymous):

\[v^2=u^2+2as\]Are you familiar with this formula?

OpenStudy (anonymous):

\[0=u^2+2*(4*10^3)*(-0.36)\]

OpenStudy (anonymous):

-0.36 is the change in position, displacement.

OpenStudy (anonymous):

\[u=\sqrt{8*10^3*0.36 }\approx55.67m/s\]

OpenStudy (anonymous):

Let's check this. \[v=u+at=-56+4*10^3t=0\]\[t=0.014\]

OpenStudy (anonymous):

\[s=ut+0.5at^2=-56(0.014)+0.5(4*10^3)(0.014)^2=-0.392\]

OpenStudy (anonymous):

Which is close enough to the real -0.36 for comfort.

OpenStudy (anonymous):

so, is that -.392 the initial speed or 55.67?

OpenStudy (anonymous):

Can someone help me out?

OpenStudy (anonymous):

why would you ask that?

OpenStudy (anonymous):

You could have indicated when you weren't getting it initially. It's the latter.